[Brahms-l] Shift report 20050220 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Feb 20 2005 - 15:55:18 EST
Shifter leader was CN
Shifters were 


1. It took 2 hours to get a new store. Otherwise good beam. First store was extended for an hour.

2. Runs
-
MRS@40A1050, FS@8B2442 -

runno(tr2, tr3) : 13330(5k,15k),13331(78k, 247k), 13332(55k, 374k)
-
r13333 in progress.

3. T1 had the problem with the data led being on even when where was no run. It was solved by having it switched off for an hour between the stores....

T5 couldnt turn on. However this was solved though correspondance with FV.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Feb 20 15:55:36 2005

This archive was generated by hypermail 2.1.8 : Sun Feb 20 2005 - 15:55:49 EST